OpenStudy (anonymous):

How do you solve ((a+1)/3)+((a+1)/4)=9 ?

OpenStudy (aayushi.somani):

i dont know the answer is correct or not..! but this ques can be solved like this.. [(a+1)/3]+[(a+1)/4]=9, \[\frac{ 4a+4+3a+3 }{ 12 }=9\], 7a+7=108 7(a+1)=108 a+1=15.42 a=14.42

OpenStudy (unklerhaukus):

that method works, (if you round correctly)

OpenStudy (anonymous):

but i'm suppose get 12

OpenStudy (aayushi.somani):

hey iam not able to catch my mistake ..!! whats it.??

OpenStudy (unklerhaukus):

\[ \frac{a+1}3+\frac{a+1}4=9\\ 4\frac{a+1}{12}+3\frac{a+1}{12}=9\\ 7(a+1)=12\times9\\ a+1=\frac{12\times9}7\\ a=\frac{12\times9}7-1\\ \quad=14.43\]

OpenStudy (aayushi.somani):

thats what i did @UnkleRhaukus